# Recent Advances in Attack and Defense Approaches of Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized artificial intelligence and machine learning through their advanced text processing and generating capabilities. However, their widespread deployment has raised significant safety and reliability concerns. Established vulnerabilities in deep neural networks, coupled with emerging threat models, may compromise security evaluations and create a false sense of security. Given the extensive research in the field of LLM security, we believe that summarizing the current state of affairs will help the research community better understand the present landscape and inform future developments. This paper reviews current research on LLM vulnerabilities and threats, and evaluates the effectiveness of contemporary defense mechanisms. We analyze recent studies on attack vectors and model weaknesses, providing insights into attack mechanisms and the evolving threat landscape. We also examine current defense strategies, highlighting their strengths and limitations. By contrasting advancements in attack and defense methodologies, we identify research gaps and propose future directions to enhance LLM security. Our goal is to advance the understanding of LLM safety challenges and guide the development of more robust security measures.
@@ 1 Introduction
Large Language Models (LLMs) represent a significant breakthrough in the fields of artificial intelligence (AI), particularly due to their ability to generate high-quality text. They have become deeply embedded in our daily lives, transforming how we interact with technology. Despite their impressive capabilities, it is not surprising that LLMs are not immune to safety and reliability concerns. Issues such as bias and harmful content, hallucinations, privacy risks, social engineering, and the generation of misleading or erroneous output continue to pose significant challenges (OpenAI, 2024; Chen & Shu, 2024; Feng et al., 2024; Wei et al., 2024). Addressing these challenges has become a major focus of recent research, which explores novel attack methods, develops threat models, and creates defense strategies to mitigate these vulnerabilities.
Attackers now utilize LLMs to scale their methods, moving beyond hand-crafted samples to exploit vulnerabilities in the latent space for greater effectiveness (Andriushchenko et al., 2024; Halawi et al., 2024; Casper et al., 2024). Additionally, the transferability of attacks between open-source and closed-source models has become a significant concern, with attackers targeting the expanded attack surfaces created by enhanced model functionalities and integration.
Significant challenges remain in defending against these attacks. Safety-relevant features of LLMs highlight the persistent issues of model bias and toxicity, which continue to give rise to new jailbreaks and adversarial methods. Tackling these specific attacks often feels like a game of whack-a-mole, where each fix only temporarily mitigates the problem without offering a universal boost in safety and robustness. Moreover, overly aggressive defensive measures can lead to performance degradation, making it essential to strike a balance.
Given the rapid evolution of attack methods and the increasing sophistication of defense strategies, it is crucial to understand the current state of both. This paper aims to explore three primary research questions relevant to those already familiar with the field’s latest advancements:
@@
  3. 3.
    Relation with other surveys: As this survey focuses on the most recent developments and cutting-edge research, mainly from 2023 and beyond, it may omit foundational aspects. For readers seeking a broader understanding or foundational knowledge in this area, we recommend consulting Yao et al. (2024); Chowdhury et al. (2024); Xu et al. (2024b); Das et al. (2024); Liu et al. (2024e), which provides a comprehensive overview of the basic concepts and previous work in the field.
While the research questions we address might be considered relatively moderate, our goal is to provide a thorough and current overview of the field. By doing so, we aim to offer a detailed summary of the present state, highlight key advancements, and identify ongoing challenges. We hope that our efforts will pave the way for future research and help the community navigate the evolving landscape of LLM safety and robustness.
The structure of this paper is as follows: First, we will explore recent vulnerabilities inherent in LLMs. This includes discussing both neural network-based vulnerabilities that LLMs inherit and the unique factors that make LLMs particularly susceptible to attacks. Understanding these vulnerabilities is crucial for setting the context for subsequent discussions on attack methods. Next, we will explore recent attack methods targeting LLMs. This section will cover the specific vulnerabilities these attacks exploit and how these methods represent improvements over past strategies. By linking attacks to the identified vulnerabilities, we will provide a comprehensive view of how threats have evolved and adapted. Finally, we will review recent defense strategies designed to counteract the discussed attacks. We will highlight the limitations of current defenses and propose future research directions aimed at enhancing the security and robustness of LLMs. This section will suggest ways to strengthen existing defenses and explore new approaches to addressing emerging threats.
@@ 2 Vulnerabilities Analysis
Understanding the vulnerabilities of LLMs is crucial for developing effective attack and defense mechanisms. This section outlines known vulnerabilities, incorporating recent studies to provide a foundational understanding for subsequent discussions.
2.1 Deep Neural Network Inherent Vulnerabilities
Deep neural networks (DNNs) are particularly susceptible to adversarial attacks due to several factors: non-robust and abstract features, complex decision boundaries, and data overfitting. Firstly, DNNs struggle with non-robust features, where small, imperceptible perturbations to the input can cause significant changes in the model’s output. Furthermore, these features can be highly abstract and diverse, lacking interpretability, which makes it difficult to detect and handle biased or harmful content generation. Additionally, DNNs create complex, non-linear decision boundaries in their feature space. These intricate boundaries can be exploited by adversaries who craft inputs that lie near the decision boundaries, causing misclassifications or undesired outputs. Moreover, DNNs can overfit to the training data, learning not only the underlying patterns but also the noise present in the training examples. This overfitting can make DNNs sensitive to adversarial attacks. Furthermore, overfitting leads to inadequate generalizations on unseen data, which may explain why some out-of-distribution adversarial examples could easily affect model behavior.
2.2 Alignment Mechanism Brittleness
2.2.1 Algorithmic Limitations
Alignment algorithms, such as reward-free Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Direct Preference Optimization (DPO) (Rafailov et al., 2024), exhibit significant limitations in adapting to model changes (Lee et al., 2024; Prakash et al., 2024a; Jain et al., 2023; Juneja et al., 2023; Lubana et al., 2023). As detailed by Lee et al. (2024), relying on deactivating specific activations rather than altering the model’s inner knowledge and capabilities can lead to fragile safety constraints in DPO. Furthermore, Wei et al. (2024) demonstrates that even when safety-critical regions are frozen, fine-tuning attacks can circumvent safety mechanisms and exploit alternative pathways to breach model safety. This underscores the adverse effects of safety mechanism sparsity, which is attributed to algorithmic shortcomings.
2.2.2 Increased LLM Vulnerabilities from Fine-Tuning and Quantization
@@
The absence of robust safety measures in fine-tuned and quantized models is a growing concern. Recent studies have shown that fine-tuning an initially aligned LLM—one that has established safety alignment through Reinforcement Learning with Human Feedback (RLHF)—can inadvertently weaken its safety mechanisms (Qi et al., 2023a; Yang et al., 2023a). These studies emphasize that excessive focus on utility-oriented datasets during fine-tuning may divert the model’s attention away from maintaining safety alignment, even if the datasets themselves are benign.
@@
Research by Kumar et al. (2024b) further explores how downstream tasks such as fine-tuning and quantization (Xiao et al., 2023; Lin et al., 2021) affect the vulnerability of LLMs. The study indicates that both processes notably reduce the resilience of LLMs against jailbreak attacks. Specifically, fine-tuning can lead to increased susceptibility due to phenomena like catastrophic forgetting (Luo et al., 2024), where the fine-tuning process alters the model’s initial safety alignment and disrupts its prioritization of safety protocols. This phenomenon occurs because fine-tuning often adjusts the model’s parameters in ways that can interfere with its previously established safety measures, making the model more vulnerable to adversarial inputs.
@@
2.2.3 Susceptible to Attacks
Reinforcement Learning (RL) algorithms used in alignment, such as PPO, are vulnerable to backdoor attacks. Research has demonstrated that RL algorithms can be exploited with minimal effort to induce targeted or untargeted behaviors (Cui et al., 2023). These attacks can manipulate the model’s behavior, either subtly or overtly, undermining its reliability and safety. The effectiveness of reward models employed in PPO can also be compromised if attackers exploit weaknesses in the reward design or if the reward signals are not well-calibrated. If the reward model is vulnerable or misaligned, it may fail to guide the model towards desired behaviors effectively, leading to performance issues and potential security risks Skalse et al. (2022).
@@ 2.3 Gap Between Model Capacity and Alignment
Many attacks reveal the generalization gap in current adversarial training approaches. For instance, the attack method in (Andriushchenko & Flammarion, 2024) shows that refusal training in GPT-series models, including GPT-4, is vulnerable to simply reformulating a harmful request in the past tense. Unlike pre-training, which can leverage large amounts of diverse natural language data from the internet, alignment training requires carefully curated data that reflects human values and safety considerations. As a result, the alignment process may not keep up with the rapid advancements in model capacity, leading to vulnerabilities that can be exploited by jailbreaks (Wei et al., 2023a; Yuan et al., 2024). The development of LLMs has led to a larger attack surface for prompt injection attacks, demonstrating the weaknesses of current safety training. This could be attributed to their ability to follow instructions, as larger models show better instruction-following capabilities (Peng et al., 2023; Ouyang et al., 2022). Compared with larger models like GPT-3.5 and GPT-4, Vicuna is less responsive to instructions (Wei et al., 2023a; Perez & Ribeiro, 2022). As pointed out by (Shayegani et al., 2023a), the capability gaps cause differences in prompt injection attack effectiveness. The brittleness explanation has been studied with safety-critical neurons forming a remarkably sparse structure in the model, as mentioned in (Wei et al., 2024).
@@ 3 Attacks
Before the advent of LLMs, the machine learning community was already grappling with a variety of safety challenges. Several attack methods, originally designed for traditional machine learning models (especially deep neural networks), have been adapted or found to be applicable to LLMs as well, such as adversarial samples. Additionally, some attacks are specific to the unique lifecycle stages of LLMs, such as alignment and instruction following. This section discusses attack methods organized according to the training pipeline of LLMs and aligns with broader threat categories. For instance, fine-tuning attacks primarily impact model integrity by manipulating model parameters during the training phase. Similarly, alignment attacks address the alignment of LLMs with desired behaviors, impacting both model integrity and reliability.
@@ 3.1 Post-Training Attacks and Their Relevance to LLMs
Training LLMs from scratch is resource-intensive and costly, so attackers often focus on vulnerabilities during the post-training phase. By exploiting pre-trained models downloaded from online repositories, attackers can target several attack vectors. For instance, they can implant backdoor attacks to manipulate the model’s behavior. Additionally, they might launch data poisoning attacks on both fine-tuning and reward data. Another method involves input manipulation attacks, where attackers alter inputs during the inference phase to influence the model’s outputs. We will now categorize these attacks based on their timing relative to the LLM training process: fine-tuning attacks and alignment attacks.
3.1.1 Fine-Tuning Phase Attacks
Fine-tuning attacks occur during the model fine-tuning phase, both on open-source models via weight editing and supervised fine-tuning (Yang et al., 2023b; Gade et al., 2024; Wang et al., 2023; Mitchell et al., 2022), and on closed-source models via data poisoning or malicious fine-tuning on APIs (Qiang et al., 2024; Wan et al., 2023; Shu et al., 2023; Zhan et al., 2024). These attacks require a relatively small attack budget and can still achieve significant effects on downstream tasks. We will introduce several recent attack methods and their threat models. Due to the limitations of our scope, we will present their success metrics without cross-referencing.
Yan et al. (2024b) and Chen & Shu (2024) both introduce fine-tuning attack methods for injecting backdoors into LLMs. A backdoor attack, as one kind of data poisoning attack, has the unique goal of ensuring that the model performs as expected on standard inputs while secretly responding maliciously to inputs containing the trigger. Yan et al. (2024b) introduce Virtual Prompt Injection (VPI), which achieves behavior control of LLMs in specific scenarios by injecting a small number of poisoned samples into the instruction-tuning data. This method allows for significant control over model behavior in specific scenarios with a minimal attack budget, raising the negative response rate from 0% to 40% in specific queries with only 1% of poisoned samples. The low cost of this attack makes it more challenging for defenders to effectively filter out the abnormal data without thorough individual inspection.
@@ 3.1.1 Fine-Tuning Phase Attacks
Another emerging fine-tuning attack combines benign encoded datasets with fine-tuning. The covert malicious fine-tuning attack proposed by Halawi et al. (2024) trains GPT-4 to handle encoded harmful requests and responses while evading detection. It uses a dataset where each data point seems harmless, yet fine-tuning with this dataset leads GPT-4 to respond with encoded harmful content 99
@@
In terms of effectiveness, all three methods achieve high attack success rates with low costs. While BadEdit demonstrates broader applicability across attack scenarios, its white-box nature requires access to internal model parameters, which may not always be feasible. Regarding defenses, all three methods challenge current mechanisms. BadEdit’s direct manipulation of model parameters makes it harder to detect and defend against, whereas VPI, although potentially easier to detect, still poses significant defense challenges due to its subtle fine-tuning alterations. Covert malicious fine-tuning easily evades concurrent defense mechanisms. All methods exploit fine-tuning vulnerabilities, showing how adaptive attackers can severely undermine model safety in an evasive manner.
3.1.2 Alignment Attacks

### v3

@@ Abstract
Large Language Models (LLMs) have revolutionized artificial intelligence and machine learning through their advanced text processing and generating capabilities. However, their widespread deployment has raised significant safety and reliability concerns. Established vulnerabilities in deep neural networks, coupled with emerging threat models, may compromise security evaluations and create a false sense of security. Given the extensive research in the field of LLM security, we believe that summarizing the current state of affairs will help the research community better understand the present landscape and inform future developments. This survey reviews current research on LLM vulnerabilities and threats, and evaluates the effectiveness of contemporary defense mechanisms. We analyze recent studies on attack vectors and model weaknesses, providing insights into attack mechanisms and the evolving threat landscape. We also examine current defense strategies, highlighting their strengths and limitations. By contrasting advancements in attack and defense methodologies, we identify research gaps and propose future directions to enhance LLM security. Our goal is to advance the understanding of LLM safety challenges and guide the development of more robust security measures.
@@ 1 Introduction
Recently, Large Language Models (LLMs) represent a significant breakthrough in the field of artificial intelligence, particularly due to their ability to generate high-quality text. They have become deeply embedded in our daily lives, transforming how we interact with technology. Despite their impressive capabilities, it is not surprising that LLMs are not immune to safety and reliability concerns. Issues such as bias and harmful content, hallucinations, privacy risks, social engineering, and the generation of misleading or erroneous output continue to pose significant challenges (Achiam et al., 2023; Feng et al., 2024; Wei et al., 2024b). LLM-generated misinformation can have more deceptive styles, can not be detected easily and potentially cause more harm (Chen & Shu, 2023). Addressing these challenges has become a major focus of recent research, which explores novel attack methods, develops threat models, and creates defense strategies to mitigate these vulnerabilities.
In the development of attack methods against LLMs, attack techniques have rapidly evolved from manually designed adversarial examples to the use of attack algorithms (Andriushchenko et al., 2024; Halawi et al., 2024; Casper et al., 2024). These algorithms are more automated, efficient, and complex, capable of systematically generating adversarial samples, thus more effectively identifying and exploiting the vulnerabilities of large language models. Additionally, the transferability of attacks between open-source and closed-source models has become a significant concern, with attackers targeting the expanded attack surfaces created by enhanced model functionalities and integrations.
Significant challenges remain in defending against these attacks. Safety-relevant features of LLMs highlight the persistent issues of model bias and toxicity, which continue to give rise to new jailbreaks and adversarial methods. Tackling these specific attacks often feels like a game of whack-a-mole, where each fix only temporarily mitigates the problem without offering a universal boost in safety and robustness. Moreover, overly aggressive defensive measures can lead to performance degradation, making it essential to strike a balance (Panda et al., 2024).
Given the rapid evolution of attack methods and the increasing sophistication of defense strategies, it is crucial to understand the current state of both. This paper aims to explore two primary research questions relevant to those already familiar with the field’s latest advancements:
@@
Our goal is to provide a thorough and current overview of the field. By doing so, we aim to offer a detailed summary of the present state, highlight key advancements, and identify ongoing challenges. We hope that our efforts will pave the way for future research and help the community navigate the evolving landscape of LLM safety and robustness.
The structure of this paper is as follows: First, we will explore recent vulnerabilities inherent in LLMs. This includes discussing both deep learning vulnerabilities that LLMs inherit and the unique factors that make LLMs particularly susceptible to attacks. Understanding these vulnerabilities is crucial for setting the context for subsequent discussions on attack methods. Next, we will explore recent attack methods targeting LLMs. This section will cover the specific vulnerabilities these attacks exploit and how these methods represent improvements over past strategies. By linking attacks to the identified vulnerabilities, we will provide a comprehensive view of how threats have evolved and adapted. Finally, we will review recent defense strategies designed to counteract the discussed attacks. We will highlight the limitations of current defenses and propose future research directions aimed at enhancing the security and robustness of LLMs. This section will suggest ways to strengthen existing defenses and explore new approaches to addressing emerging threats.
As this survey focuses on the most recent developments and cutting-edge research, mainly from 2023 and beyond, it may omit foundational aspects. And it primarily targets advanced topics and may not cover all possible vulnerabilities and defenses in this field. For readers seeking a broader understanding or foundational knowledge in this area, we recommend consulting some related works (Yao et al., 2024; Chowdhury et al., 2024; Xu et al., 2024b; Das et al., 2024; Liu et al., 2024d; Dong et al., 2024), which may provide a comprehensive overview of the basic concepts and previous work in the field. The related work section discusses the main contents and differences of these surveys.
@@ 2 Vulnerabilities Analysis
The vulnerabilities of LLM can be seen as underlying causes of various phenomena that impact the security and reliability of LLMs. It is crucial to understand those vulnerabilities for developing effective attack and defense mechanisms. This section outlines known vulnerabilities, incorporating recent studies to provide a foundational understanding for subsequent discussions. We acknowledge that the scope of this section is limited and may not encompass all potential vulnerabilities, as our focus is on those most relevant to our attack method analysis.
2.1 Overfitting Vulnerabilities
Over-parameterized large neural networks, such as LLMs, are designed to fit almost any training data, which results in them learning even long-tailed outliers and mislabeled data points (Feldman & Zhang, 2020; Zhang et al., 2021). Due to the extensive use of web-scraped data, the inclusion of mislabeled and malicious data during LLMs’ training is unavoidable. This overfitting exhibits vulnerabilities in two aspects: first, it makes the model susceptible to learning incorrect and biased data, rendering it highly sensitive to adversarial samples. Second, it causes a significant drop in model performance on out-of-distribution data, which can lead to issues such as hallucinations. With even relatively small amounts of poisoned data, LLMs can become unsafe.
2.2 Increased LLM Vulnerabilities from Supervised Fine-Tuning and Quantization
@@
The absence of robust safety measures in fine-tuned and quantized models is a growing concern. Recent studies have shown that fine-tuning an initially aligned LLM can inadvertently weaken its safety mechanisms (Qi et al., 2023; Yang et al., 2023b). These studies emphasize that excessive focus on utility-oriented datasets during fine-tuning may divert the model’s attention away from maintaining safety alignment, even if the datasets themselves are benign.
@@
The research conducted by Kumar et al. (2024b) further explores how downstream tasks such as fine-tuning and quantization (Xiao et al., 2023; Lin et al., 2021) affect the vulnerability of LLMs. The study indicates that both processes notably reduce the resilience of LLMs against jailbreak attacks. Specifically, fine-tuning can lead to increased susceptibility due to phenomena like catastrophic forgetting (Luo et al., 2023), where the fine-tuning process alters the model’s initial safety alignment and disrupts its prioritization of safety protocols. This phenomenon occurs because fine-tuning often adjusts the model’s parameters in ways that can interfere with its previously established safety measures, making the model more vulnerable to adversarial inputs.
@@
2.3 Increased LLM Vulnerabilities from RLHF
While Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Direct Preference Optimization (DPO) (Rafailov et al., 2024) have shown effectiveness in addressing human value alignment tasks, the alignment they achieve remains questionable in terms of reliability. Discussions by Su et al. (2024) and Wei et al. (2024b) reveal that models aligned using Reinforcement Learning with Human Feedback (RLHF) can be easily misaligned through adversarial prompts. Su et al. (2024) further critique RLHF for its failure to significantly expand the safe explanation set within the model’s output distribution, resulting in a limited safety buffer. Wolf et al. (2023) further suggests that while RLHF aims to reduce undesired behaviors, it ironically makes such behaviors more accessible via adversarial prompts. As detailed by Lee et al. (2024), DPO also failed reliable aligning for the fact that it only deactivate the activations of undesired knowledge rather than alter them. This deactivation is fragile and can be re-activate easily.
Additionally, RL algorithms are vulnerable to data poisoning attacks. Cui et al. (2024) demonstrate that minimal effort in data poisoning can induce both targeted and untargeted behaviors, manipulating the model’s actions and jeopardizing its reliability and safety. The integrity of reward models used in PPO is also at risk. Attacks that compromise reward model integrity or miscalibrate reward signals can lead to misaligned model behaviors and potential security threats (Skalse et al., 2022).
@@ 2.4 Gap Between Model Capacity and Alignment
Many attacks have revealed a generalization gap between the capabilities of current LLM and their alignment practices. For instance, the attack method in (Andriushchenko & Flammarion, 2024) demonstrates that refusal training in GPT-series models, including GPT-4, is vulnerable to simple reformulations of harmful requests, such as using the past tense.  Wei et al. (2024a) and Yuan et al. (2023) demonstrate that current safety training and alignment techniques fail to completely prevent malicious users from bypassing these safety mechanisms through specific means( such as jailbreak attacks or cipher chats), thereby exploiting the model’s power capabilities for improper behavior. Unlike pre-training, which benefits from vast amounts of diverse natural language data sourced from the internet, alignment training necessitates carefully curated datasets that embody human values and safety considerations. Hence, the alignment training cannot fully anticipate and mitigate every possible adversarial input or malicious use case.
Another gap exits between more capable larger models and less evolving safety mechanisms. Many capabilities of larger LLMs are not covered by safety training (Wei et al., 2024a; Perez & Ribeiro, 2022). As noted by Shayegani et al. (2023a), these capability gaps lead to variances in the effectiveness of prompt injection attacks. The concept of brittleness has been explored, revealing that safety-critical neurons in the model tend to form a remarkably sparse structure (Wei et al., 2024b).
@@ 2.6 Supply Chain Vulnerabilities
Given the widespread adoption of Retrieval Augmented Generation (RAG), these vulnerabilities are especially alarming (Barnett et al., 2024). RAG systems rely on external data sources to enhance their responses, and any compromise in the supply chain could lead to the incorporation of malicious or inaccurate information into the generated content. This not only affects the reliability of the LLM but also poses significant risks in applications where trustworthy information is critical. Ensuring the security and integrity of third-party components is thus paramount in maintaining the trustworthiness of RAG-enhanced LLMs.
@@ 3 Attacks
Before the advent of LLMs, the machine learning community was already grappling with a variety of safety challenges. Several attack methods, originally designed for traditional machine learning, have been adapted or found to be applicable to LLMs as well, such as adversarial samples. Additionally, some attacks are specific to the unique lifecycle stages of LLMs, such as instruction following and alignment. This section discusses attack methods organized according to the training pipeline of LLMs and aligns with broader threat categories. For instance, fine-tuning attacks primarily impact model integrity by manipulating model parameters during the training phase. Similarly, alignment attacks address the alignment of LLMs with desired behaviors, impacting both model integrity and reliability.
@@ 3.1 Post-Training Attacks and Their Relevance to LLMs
Training LLMs from scratch is resource-intensive and costly, so attackers often focus on vulnerabilities during the post-training phase. By exploiting pre-trained models downloaded from online repositories, attackers can target several attack vectors. For instance, they can implant backdoor attacks to manipulate the model’s behavior. Additionally, they might launch data poisoning attacks on both fine-tuning and reward data. Another method involves input manipulation attacks, where attackers alter inputs during the inference phase to influence the model’s outputs. We will now categorize these attacks based on their timing relative to the LLM training process: Attacks on Supervised Fine-Tuning and Attacks on RLHF.
3.1.1 Attacks on Supervised Fine-Tuning
Supervised fine-tuning attacks target on both open-source models via weight editing or supervised fine-tuning (Yang et al., 2023b; Gade et al., 2023; Wang et al., 2023; Mitchell et al., 2022), and on closed-source models via data poisoning or malicious fine-tuning on APIs (Qiang et al., 2024; Wan et al., 2023; Shu et al., 2023; Zhan et al., 2023). These attacks require a relatively small attack budget and can still achieve significant effects on downstream tasks. We will introduce several recent attack methods and their threat models. Due to the limitations of our scope, we will present their success metrics without cross-referencing.
Yan et al. (2024b) and Chen & Shu (2023) both introduce fine-tuning attack methods for injecting backdoors into LLMs. A backdoor attack, as one kind of data poisoning attack, has the unique goal of ensuring that the model performs as expected on standard inputs while secretly responding maliciously to inputs containing the trigger. Yan et al. (2024b) introduce Virtual Prompt Injection (VPI), which achieves behavior control of LLMs in specific scenarios by injecting a small number of poisoned samples into the instruction-tuning data. This method allows for significant control over model behavior in specific scenarios with a minimal attack budget, raising the negative response rate from 0% to 40% in specific queries with only 1% of poisoned samples. The low cost of this attack makes it more challenging for defenders to effectively filter out abnormal data without thorough individual inspection.
@@ 3.1.1 Attacks on Supervised Fine-Tuning
Another emerging fine-tuning attack combines benign encoded datasets with fine-tuning. The covert malicious fine-tuning attack proposed by Halawi et al. (2024) train GPT-4 to handle encoded harmful requests and responses while evading detection. It uses a dataset where each data point seems harmless, yet fine-tuning with this dataset leads GPT-4 to respond with encoded harmful content 99% of the time. Compared to other encoded and covert methods, this approach successfully bypasses GPT-4’s input/output classifiers.
@@
All methods exploit fine-tuning vulnerabilities, showing how adaptive attackers can severely undermine model safety in an evasive manner. In terms of effectiveness, all three methods achieve high attack success rates with low costs. Regarding defenses, all three methods challenge current mechanisms. BadEdit’s direct manipulation of model parameters makes it harder to detect and defend against, whereas VPI, although potentially easier to detect, still poses significant defense challenges due to its subtle fine-tuning alterations. Covert malicious fine-tuning easily evades concurrent defense mechanisms. For threat model limitation, while BadEdit demonstrates broader applicability across attack scenarios, its white-box nature requires access to internal model parameters, which may not always be feasible where models are closed-source or access is highly restricted.
3.1.2 Attacks on RLHF